# Ring Deconvolution Microscopy: An Exact Solution for Spatially-Varying Aberration Correction

### v3

@@ 2.4 High NA multicolor fluorescence microscopy
After calibration, we image BPAE cells and process them with both RDM and standard deconvolution. Figure 3 b) shows that RDM deblurs the raw images consistently over the entire FoV, including the corners of the image, while standard deconvolution becomes low contrast and noisy near the edges. In both examples, RDM is able to resolve subcellular features in the actin and mitochondria near the edges that are not visible in standard deconvolution. Such capability allows for larger FoVs to be used, lessening the burden of stitching together many smaller FoV images when the sample is large.
@@
Figure 4: RDM for point-scanning micro-endoscopy through a multimode fiber. Comparison of deconvolution (middle-left), spatially-varying Richardson-Lucy or SVRL (middle-right), and RDM (right) on images taken from a point-scannning multimode fiber micro-endoscope [15]. Top row: 1 $\mu m$ beads; all methods resolve the circularly-shaped beads in the center (blue inset). Away from the center, deconvolution and SVRL fail to completely remove the ellipiticity of the bead but RDM does (green inset). Moreover, in the corner only RDM can resolve bead clusters into their component beads (yellow inset). Bottom row: live Wistar rat neuron; RDM can clearly resolve the 2-shaped structure (blue inset), sharpen the neuron spine (green inset), and reveal a point-like structure (yellow inset) near the edges where the other methods cannot.
2.5 Point-scanning micro-endoscopy through a multimode fiber
@@
Point-scanning micro-endoscopy through a multimode fiber as in [15, 38] is a powerful technique for deep in vivo imaging at subcellular resolution, with particular applications in the brain and other sensitive organs, where minimal tissue damage is required. However, due to the extreme constraints imposed in the design of the fiber, its resolution capabilities degrade quickly and severely away from the center of the image, resulting in a smaller usable FoV (see top row of Fig. 4). The spatially-varying images from such a system have been heuristically deblurred in [15], but RDM, with its theoretically exact formulation, offers a significant improvement in performance with far less calibration.
@@
To verify this, we process images of beads and live rat neurons from [15] (refer to their work for additional details) using RDM, their spatially-varying Richardson-Lucy (SVRL), and standard deconvolution, for comparison. SVRL is calibrated by measuring a uniform grid of 21 by 21 PSFs—441 images—whereas, RDM is calibrated with a single image of a few randomly placed beads. Despite the considerably lighter calibration and similar computational complexity, RDM provides a significant image quality improvement over SVRL as demonstrated in Fig. 4. In the corners of the bead image, we see that RDM is able to remove the aberration-induced ellipticity of the underlying circular shaped beads and resolve clumps of beads, unlike the other methods. The same holds for the neuron images—in the corners, RDM is able to resolve distinctive subcellular features and tighten the spread of thin spines far better than the other methods.
@@
Figure 5: Simulations to quantify RDM performance. a) Error maps show the normalized absolute difference between the blur produced by deconvolution/ring deconvolution and the “true” blur (produced by manually superimposing every PSF at every pixel). When off-axis aberrations are small (top row), both forward models are accurate. When aberrations are large (bottom row), convolution becomes significantly worse, yet ring convolution remains accurate. b) Forward model mean squared error (MSE) as a function of off-axis aberration magnitude (run on the image from a)). c) Runtime of each method as a function of size of the image (in pixels); computing the true blur quickly becomes intractable, but convolution and ring convolution remain relatively fast. Times averaged over n=50 trials. d) A noisy image of randomly placed PSFs is used to estimate the underlying Seidel coefficients, which can then be used to generate PSFs at any location. e) Seidel fitting error is plotted as a function of iteration in the optimization algorithm to demonstrate convergence. Each purple line is a different trial of n=50 trials with a randomly sampled set of underlying Seidel coefficients. The red dashed line is the per-iteration median. f) Mean squared error of the fitted Seidel coefficients plotted as a function of SNR of the calibration image. The average coefficient error is plotted along with the variance (error bar) over n=50 random trials. Some example calibration PSFs are shown. Even under severe noise, the Seidel fit is still accurate. g) Deblurring results on noisy images from the CARE dataset, with PSNR values above each method. h) Zoom-ins of an off-axis patch in each deblurred image; ring deconvolution (D) and DeepRD (E) have the highest quality. i) Accuracy (PSNR) vs runtime of each method (averaged over n=28 true blurred images using unseen coefficients), with the number of model parameters (written below each circle) determining the size of the circle.
2.6 Quantifying performance improvements
To quantify the performance of our methods, we conducted a series of simulations in which we have access to the unblurred ground-truth image. First, we verify the accuracy of ring convolution as a forward model. Second, we quantify the accuracy of fitting Seidel coefficients to a single calibration image, even in high-noise conditions. Finally, we demonstrate the effectiveness of ring deconvolution and DeepRD in deblurring images from a spatially-varying system.
2.6.1 Forward model: Ring convolution
Simulating the blurring operation of an imaging system is a critical part of image deblurring. By mathematically specifying the forward model of the system, we can then invert it in order to obtain the desired deblurring algorithm. When a system is space invariant, the forward model is a simple convolution operation. For space-varying systems, however, we must account for the changes in the PSF across the FoV. The brute force approach for doing so would superimpose the PSFs at each pixel in the image to compute the ’true blur’, at the cost of long compute times. However, when the system is rotationally symmetric (i.e., varies only radially), ring convolution is an equivalent operation to the brute force method, but runs much quicker—on the order of seconds—even for image sizes upward of $512\times 512$.
@@ 2.6.1 Forward model: Ring convolution
To verify, we blur each test image using spatially-varying PSFs rendered from a randomly chosen set of Seidel coefficients. We treat the ’true blur’ as ground truth and compare the error maps for both traditional LSI convolution and our ring convolution. As expected, traditional convolution results in errors near the edges of the FoV, whereas ring convolution produces accurate blur across the entire image (Fig. 5a). We can quantify the severity of the spatial variance by calculating the norm of the off-axis Seidel coefficients; in Fig. 5b we see that the error for traditional convolution increases approximately linearly with this aberration magnitude. In contrast, ring convolution continues to produce an accurate blur, independent of the strength of the aberrations. In Fig. 5c, we compare compute times for these forward models, showing that our ring convolution method is significantly faster than the other exact method (true blur).
2.6.2 Seidel fitting
@@
Next, we verify our ability to accurately fit Seidel coefficients from a single noisy image of randomly scattered point sources. As detailed in Sec. 4.2, the Seidel fitting procedure involves searching for the set of 5 primary Seidel coefficients that best fit the measured PSFs at their given positions in the calibration image. We simulate this process many times by randomly generating sets of Seidel coefficients, using them to produce a calibration image with additive Gaussian noise, and then estimating those coefficients only using the noisy calibration image (see Fig. 5d). Section 4.2 shows that this optimization problem is nonconvex, yet we find that our estimated Seidel coefficients almost always converge correctly. This is not entirely surprising since the Seidel procedure is optimal in the maximum likelihood sense [39]. To quantify, we ran many Seidel fits over random coefficients, randomly located PSFs, and random draws of additive noise, and plotted the error between estimated and true coefficients in Fig. 5e. Due to the nonconvexity, we see that not every case produces errors which converge to 0. However, two things provide assurance: 1)the median convergence approaches 0, meaning that a majority of optimizations will produce the optimal solution, and 2) even the runs which do not converge to the global minimum produce PSFs which are close enough to the true ones to provide good quality ring deconvolution .
Finally, since the Seidel fitting procedure acts as a PSF denoiser, we tested the fit with varying amounts of noise (up to $-20$ dB SNR) and found that the fit only reduces performance slightly, even with severe additive noise (Fig. 5f).
2.6.3 Spatially-varying deblurring
@@
After verifying that our forward model and Seidel fitting methods perform well, we next use them in an inverse problem to deblur images and quantify the performance. We compare our methods (Ring deconvolution and DeepRD) with traditional deconvolution and the baseline CARE-based U-Net. The learning methods are trained with images from the CARE and Div2k datasets [21, 37], after synthetically blurring them with space-varying PSFs rendered from a random set of Seidel coefficients. It was computationally infeasible to generate the training set using the true blur technique, so instead we used our ring convolution to generate the blurred images for training. This should not pose a major issue since ring convolution is theoretically equivalent to the true blur case. To remove any doubt, we construct the test set using the true blur. Each model is first pretrained on 80,500 data points from the Div2k dataset (800 base images blurred with 100 different Seidel coefficients) and then fine-tuned on a small 8400 data point subset of the CARE dataset (24 images, 350 different Seidel coefficients). After training, each method is tested on a test of 28 unseen images from the CARE dataset, which are blurred using the true blur method and noised with additive Gaussian noise.
The results of each method on one representative test image are shown in Fig. 5g,h, with the PSNR (in dB) listed above. Both DeepRD and ring deconvolution deblur better near the edges and corners of the image, where the PSF most deviates from the center PSF. Despite using ring convolution for the training set and the true blur for the test set, neither of the networks (U-Net and DeepRD) show signs of model mismatch.
@@
In Fig. 5i, we plot the average accuracy (PSNR) vs runtime across the full test dataset, with the size of the circle representing the number of parameters needed (memory footprint) for each method. Ring deconvolution provides the best reconstruction, but it is also the largest and slowest of the methods tested (note that inverting the full blur would take hundreds of hours per image and is consequently not included). DeepRD performs nearly as well and has the fewest needed parameters of all the space-varying techniques, allowing it to be fast and memory efficient. The baseline U-Net and standard deconvolution PSNR values are significantly worse.
Figure 6: RDM for shift-invariant systems. We run our Seidel and blind deconvolution (left of gray line) on raw images (leftmost column) of tardigrades (bottom two rows) and rabbit liver tissue (top row). We also run standard deconvolution (2nd from the left) for comparison. Our methods outperform standard deconvolution by using a synthetic PSF, which prevents artifacts and loss of resolution from noise-based artifacts. In the first row, our Seidel deconvolution is able to resolve elements 5 and 6 of group 9 in the USAF resolution chart, which are not resolvable by the other methods. In the second row, both Seidel and blind deconvolution reduce ringing artifacts present in standard deconvolution. Our blind methods usually matches the non-blind Seidel deconvolution since it usually estimates the same spherical aberration coefficient; however, as in row 1, it can overestimate the coefficient resulting in an overly sharp image.

### v5

@@ 2.2 Experimental Results
After calibration, we image BPAE cells and process them with both RDM and standard deconvolution, for comparison. Figure 3 b) shows that RDM deblurs the raw images consistently over the entire FoV, including the corners of the image, while standard deconvolution becomes low contrast and noisy near the edges. In both examples, RDM is able to resolve subcellular features in the actin and mitochondria near the edges that are not visible in standard deconvolution. Such capability allows for larger FoVs to be used, lessening the burden of mechanically scanning and stitching together many smaller FoV images when the sample is large.
Figure 3: RDM on multicolor fluorescence imaging. a) Fluorescent beads imaged with a 100x 1.4 NA objective and b) corresponding Seidel-fitted PSFs demonstrate spatially-varying nature of the system. c) Two examples of Bovine Pulmonary Artery Endothelial (BPAE) cells processed by standard deconvolution and RDM. Deconvolution and RDM perform similarly in the center but RDM is better in the corner, revealing submicron features in the actin (orange arrow). RDM similarly resolves actin filaments and mitochondria where deconvolution does not.
@@
Figure 4: RDM for point-scanning micro-endoscopy through a multimode fiber. Comparison of deconvolution, spatially-varying Richardson-Lucy (SVRL), and RDM on images taken from a point-scannning multimode fiber micro-endoscope [Turcotte:20]. SVRL results are directly obtained from [Turcotte:20]. Top row: Sample with 1 $\mu m$ beads; all methods resolve the circularly-shaped beads in the center (blue inset). Away from the center, however, deconvolution and SVRL fail to completely remove the ellipiticity of the bead but RDM does (green inset). Moreover, in the corner only RDM can resolve bead clusters into their component beads (yellow inset). Bottom row: live Wistar rat neuron; RDM can clearly resolve structures (blue inset), sharpen the neuron spine (green inset), and reveal a point-like structure (yellow inset) near the edges where the other methods cannot.
Micro-endoscopy through a multimode fiber
@@
Point-scanning micro-endoscopy through a multimode fiber [Turcotte:20, vasquez2018subcellular] is a powerful technique for deep in vivo imaging at subcellular resolution, with applications in the brain and other sensitive organs, where minimal tissue damage is required. However, due to the extreme constraints imposed in the design of the fiber, its resolution capabilities degrade rapidly and severely away from the center of the image, resulting in a small usable FoV (see the top row of Fig. 4). The spatially-varying images from such a system have been heuristically deblurred in [Turcotte:20], but RDM, with its rigorous formulation, offers improvement in performance with far less calibration.
@@
To verify this, we process images of beads and live rat neurons from a previous paper [Turcotte:20] using RDM, their spatially-varying Richardson-Lucy (SVRL) algorithm, and standard deconvolution, for comparison. SVRL is similar to the modal decomposition work in [yanny2020miniscope3d, flicker2005anisoplanatic]. It uses fewer synthetic PSFs than RDM ( 30 vs.  120), so requires slightly less memory, but requires significantly more calibration images. Here, SVRL is calibrated by measuring a uniform grid of 21$\times$21 PSFs—441 images—whereas RDM is calibrated with a single image of a few randomly-placed beads. Despite the considerably lighter calibration and similar computational complexity, RDM provides an improvement in image quality over SVRL (see Fig. 4). In the corners of the bead image, we see that RDM is able to remove the aberration-induced ellipticity of the underlying circular beads and resolve clumps of beads, unlike the other methods. The same holds for the neuron images—in the corners, RDM is able to resolve distinctive subcellular features and tighten the spread of thin spines far better than the other methods. Additional experiments on this system and comparisons with more methods can be found in Appendix A.
@@
Figure 5: Simulations to quantify RDM performance. a) Forward model: ring convolution is compared with standard deconvolution by computing error maps that measure the absolute difference between the simulated blurs from each method and the ‘true blur’(produced by manually superimposing every PSF at every pixel). When off-axis aberrations are small (top row), both forward models are accurate. When aberrations are large (bottom row), convolution becomes significantly worse, yet ring convolution remains accurate. b) Forward model mean-squared error (MSE) as a function of off-axis aberration magnitude (for the image above). c) Runtime of each method as a function of size of the image (in pixels), averaged over n=50 trials. Computing the true blur quickly becomes intractable, but convolution and ring convolution remain relatively fast. d) A noisy image of randomly-placed PSFs is used to estimate the underlying Seidel coefficients, which can then be used to generate PSFs at any location. e) Seidel fitting error is plotted as a function of iteration in the optimization algorithm to demonstrate convergence. Each purple line is a different trial of n=50 trials with a randomly-sampled set of underlying Seidel coefficients. The red dashed line is the per-iteration median. f) Average MSE of the fitted Seidel coefficients plotted as a function of SNR of the calibration image (with standard deviation shown by error bars) over n=50 random trials. Some example calibration PSFs are shown. Even under severe noise, the Seidel fit is still accurate. g) Deblurring results on noisy images from the CARE dataset, with PSNR values above each method. h) Zoom-ins of an off-axis patch in each deblurred image; ring deconvolution and DeepRD have the highest quality. i) Accuracy (PSNR) vs runtime of each method (averaged over n=28 true blurred images using unseen coefficients), with the number of model parameters (written below each circle) determining the size of the circle.
In summary, we find that ring deconvolution consistently produces the best reconstructions among the methods we tested on our diverse panel of imaging systems. We believe the improvement arises due to the lack of approximations and heuristics in the method’s derivation. Moreover, as compared to deep-learning-based methods, ring deconvolution undergoes no learning procedure, and thus does not transmit bias from the training data into future reconstructions. DeepRD performs similarly to ring deconvolution, but it is faster and less consistent. Both perform better than the U-Net and standard deconvolution.
2.3 Simulation Results
To quantify the performance of our methods, we conducted a series of simulations in which we have access to the unblurred ground-truth image. These simulations are done with Gaussian noise and are repeated with Poisson noise in Appendix C.
@@ 2.3 Simulation Results
Our first step is to quantify the forward model: ring convolution. Simulating the blurring operation of an imaging system is a critical part of image deblurring. By mathematically specifying the forward model of the system, we can then invert it to obtain the desired deblurring algorithm. When a system is space invariant, the forward model is a simple convolution operation. For space-varying systems, however, we must account for the changes in the PSF across the FoV. The brute-force approach for doing so would superimpose weighted PSFs at each pixel in the image to compute the ‘true blur’, at the cost of long compute times. However, when the system is rotationally symmetric (i.e., varies only radially), ring convolution is an equivalent operation to the brute-force method, but runs much quicker—less than a second—even for image sizes upward of $512\times 512$. To verify, we blur each test image using spatially-varying PSFs rendered from a randomly chosen set of Seidel coefficients. We treat the true blur as ground truth and compare the error maps for both standard LSI convolution and our ring convolution. As expected, standard convolution results in errors near the edges of the FoV, whereas ring convolution produces accurate blur across the entire image (Fig. 5a). We can quantify the aberration magnitude by calculating the norm of the off-axis Seidel coefficients; in Fig. 5b we see that the error for standard convolution increases approximately linearly with this aberration magnitude. In contrast, ring convolution continues to produce an accurate blur, independent of the strength of the aberrations. In Fig. 5c, we compare compute times for these forward models, showing that our ring convolution method is nearly four orders of magnitude faster than the other exact method (true blur) for a megapixel-sized image.
@@
Next, we verify our Seidel fitting method by quantifying the error in our estimated Seidel coefficients that were fitted from a single noisy image of randomly-scattered point sources. As detailed in Sec. 4.2, the Seidel fitting procedure involves searching for the set of 5 primary Seidel coefficients that best fit the measured PSFs at their given positions in the calibration image. We simulate this process many times by randomly generating sets of Seidel coefficients, using them to produce a calibration image with additive Gaussian noise, and then estimating those coefficients only using the noisy calibration image (see Fig. 5d). Section 4.2 shows that this optimization problem is nonconvex, yet we find that our estimated Seidel coefficients almost always converge correctly. This is not entirely surprising, since the Seidel procedure is optimal in the maximum likelihood sense [Southwell:77]. To quantify, we ran many Seidel fits over random coefficients, randomly located PSFs, and random draws of additive noise, and plotted the error between estimated and true coefficients in Fig. 5e. Due to nonconvexity, we see that not every case produces errors which converge to 0. However, two things provide assurance: 1) the median convergence approaches 0, meaning that a majority of optimizations will produce the optimal solution, and 2) even the runs which do not converge to the global minimum produce PSFs which are close enough to the true ones to provide good quality ring deconvolution. Finally, since the Seidel fitting procedure acts as a PSF denoiser, we tested the fit with varying amounts of noise (up to $-20$ dB SNR) and found that the fit only reduces performance slightly, even with severe additive noise (Fig. 5f). Additional simulations for higher-order Seidel coefficients can be found in Appendix D.1.
@@
After verifying that our forward model and Seidel fitting methods perform well, we use them as part of an inverse problem to deblur images and quantify the performance. We compare our methods (Ring deconvolution and DeepRD) with standard deconvolution and the baseline CARE-based U-Net. The learning methods are trained with images from the CARE and Div2k datasets [weigert2018content, agustsson2017ntire], after synthetically blurring them with space-varying PSFs rendered from a random set of Seidel coefficients. It was computationally infeasible to generate the training set using the true blur technique, so instead we used our ring convolution to generate the blurred images for training. This should not pose a major issue since ring convolution is theoretically equivalent to the true blur case. To remove any doubt, we construct the test set using the true blur. Each model is first pretrained on 80,500 data points from the Div2k dataset (800 base images blurred with 100 different Seidel coefficients) and then fine-tuned on a small 8400 data point subset of the CARE dataset (24 images, 350 different Seidel coefficients). After training, each method is tested on a test set of 28 unseen images from the CARE dataset, which are blurred using the true blur method and noised with additive Gaussian noise. The results of each method on one representative test image are shown in Fig. 5g,h, with the PSNR (in dB) listed above. Both DeepRD and ring deconvolution deblur better near the edges and corners of the image, where the PSF most deviates from the center PSF. Despite using ring convolution for the training set and the true blur for the test set, neither of the networks (U-Net and DeepRD) show signs of model mismatch.
@@
In Fig. 5i, we plot the average accuracy (PSNR) versus runtime across the full test dataset, with the size of the circle representing the number of parameters needed (memory footprint) for each method. Ring deconvolution provides the best reconstruction, but it is also the largest and slowest of the methods tested, taking about 60 seconds for a full image. However, if we were to try to deblur the image rigorously without using RDM, it would take hundreds of hours, despite being theoretically equivalent to ring deconvolution. Thus, ring deconvolution allows for relatively fast, accurate deblurring where it was once infeasible. DeepRD performs nearly as well as ring deconvolution and has the fewest parameters needed of all the space-varying techniques, allowing it to be fast and memory efficient. DeepRD is almost three orders of magnitude faster than RDM. The baseline U-Net and standard deconvolution PSNR values are significantly worse.
2.4 Sheet Deconvolution: Extension to Lateral Symmetry
Figure 6: Sheet deconvolution exploits lateral symmetry for light-sheet microscopy. We run sheet deconvolution on samples imaged by a light-sheet fluorescence microscope with a detection NA of 1.1 and a light-sheet NA of 0.31. a) The light-sheet microscope illumination spreads along its propagation axis (orange arrow) such that only a center strip in the FoV (marked ‘useable FoV’) is optically sectioned and hence well resolved. b) Results comparing standard 3D deconvolution and sheet deconvolution on fluorescent beads (top row) and SU8686 cells (bottom row), with maximum intensity projections along each axis. Unlike standard deconvolution, sheet deconvolution resolves features nearly as well at the edges of the FoV as in the center (see insets). c) Plots of lateral and axial FWHM of reconstructed beads (from b) as a function of their distance from the center along the spatially-varying axis. For sheet deconvolution, the FWHM values grow less in lateral and axial diameter than that of standard deconvolution and differs by at most $500$ nm in diameter from center to the top/bottom of the FoV.
Revolution-invariance is not the only form of symmetry found in practical, spatially-varying systems. For example, light-sheet fluorescence microscopy (LSFM) exhibits another type of symmetry, in which the light-sheet illumination causes spatial variance, but only along one axis (see Fig. 6a)). The thinnest section, or waist, of the light-sheet is focused in the center of the FoV and becomes thicker as a function of the distance from the center. Since the light-sheet is constant along the axis orthogonal to the focusing direction, this variance only occurs along one dimension. The resulting PSF grows in axial extent as it moves along one of the lateral dimensions, but stays constant in the other lateral dimension.
This symmetry can be used to obtain a fast spatially-varying forward model of the LSFM system. By employing the same mathematical strategy used in the ring convolution theorem, we can derive sheet convolution, an exact forward model for LSFM (assuming a space-invariant objective lens). Just like ring convolution, sheet convolution leverages symmetry by convolving over the two dimensions that are spatially-invariant while integrating over the one that does vary. As a result, sheet convolution enjoys a similar improvement in computational performance as ring convolution; instead of the $O(N^{6})$ scaling of the general three-dimensional spatially-varying forward model, it achieves a $O(N^{4}log(N))$ scaling. This improvement enables rigorous, spatially-varying LSFM deconvolution where it once was computationally infeasible. All that remains is to obtain the 3D PSF at each point along the light-sheet focusing axis. This can be done in multiple ways using only a single calibration volume; the experimental PSFs from this volume are either interpolated at the missing locations or are used to fit the parameters of a PSF model such as the Gibson-Lanni model [gibson1991experimental, li2017fast]. We tried both and found that the interpolation method is better for sparser samples, while the fitted model is better for dense samples. Details along with the derivation of sheet convolution/deconvolution are in Sec. 4.4 and simulations are in Appendix B.
To combat the inherent spatial-variance of LSFM, practitioners typically employ tiling LSFM [gao2015extend], in which only the thin lateral slab of the image that lies within the light-sheet waist is kept. To acquire the full FoV, the sample is shifted after each acquisition to ensure that each part of it is imaged within the light-sheet waist. The final image is then stitched together from multiple acquisitions. However, by using our sheet deconvolution approach, the parts of each acquisition that lie outside of the waist can be recovered computationally, thereby expanding the usable FoV of the LSFM system and reducing the number of acquisitions needed for an object with a large lateral extent. While there are techniques, like axially-swept LSFM [dean2022isotropic], which speed up the process in hardware, a purely computational solution is desirable for its ease of use, reduced irradiation, and flexibility.
Figure 6 shows a demonstration of sheet deconvolution on LSFM data. For the experiment, a 3D stack of randomly-scattered fluorescent nanospheres in agarose gel was imaged and used to obtain PSFs at each position along the spatially-varying axis (see Sec. 4.4). Then, using these PSFs, sheet deconvolution is applied to samples of two types: fluorescent beads and SU8686 cells. Unlike standard deconvolution, sheet deconvolution is able to better resolve features that do not lie in the light-sheet waist and create a more homogeneous axial resolution across the entire FoV than standard deconvolution does. We quantified the resolution improvement by measuring the lateral and axial full-width half-maximum (FWHM) of fluorescence beads across the entire FoV after applying the two deconvolution methods. As expected, the lateral FWHM does not change significantly since the PSF spatial variance occurs primarily in $z$, but the axial FWHM values from sheet deconvolution are on average 300300~{}300300 nm smaller than those from standard deconvolution with 200200~{}200200 nm less standard deviation. This increase in resolution outside of the beam waist can reveal subcellular-scale features that were previously only accessible by shifting the sample. In this experiment, sheet deconvolution ran on a volume size of $512\times 512\times 160$ in about 7 minutes, which is orders of magnitude faster than solving a spatially-varying deblurring method without leveraging symmetry. With sheet deconvolution, users of LSFM can speed up the capture of large samples while still getting high resolution across the entire FoV.